x-y-z=1
2x+y+2z=4
x+y+3z=5
solve for x,y,z

Answer by Alan3354(30993)

x-y-z=1
2x+y+2z=4
x+y+3z=5
-----------
x-y-z=1
2x+y+2z=4
------------ Add
3x + z = 5 Eqn A
---------
x-y-z=1
x+y+3z=5
--------- Add
2x + 2z = 6
x + z = 3
3x+ z = 5 Eqn A
--------- Subtract
-2x = -2
x = 1
-------
x + z = 3
z = 2
-------
x-y-z=1
1-y-2 = 1
y = -2
(1,-2,2)